i need to know how yo multiply -3 1/4 x 2 and 4/9 x 3/5
Answer by stanbon(75887)

i need to know how you multiply
-3 1/4 x 2
Change from mixed numbers to improper fractions:
= (-(4*3+1))/4 x 2
= (-(13/4)) x 2
= (-13/4) x 2
Cancel the common factor of "2" to get:
= -13/2
Convert to a mixed number:
= -6 1/2
=============================

and 4/9 x 3/5
= (4/9)(3/5)
Cancel the common factor of "3" to get:
= (4/3)(1/5)
Multiply to get:
= 4/15
===============================
Cheers,
Stan H.